OpenStudy (anonymous):

Function even, odd, or neither. 1. F(x)=2x/|x|

OpenStudy (freckles):

enter in -x

OpenStudy (freckles):

if F(-x)=F(x) then F is even if F(-x)=-F(x) then F is odd if you get neither of those if parts then F is neither odd or even

OpenStudy (freckles):

\[F(-x)=\frac{2(-x)}{|-x|}=\frac{2(-x)}{|-1 \cdot x|}=\frac{-2x}{|-1| \cdot |x|}=?\]

OpenStudy (anonymous):

even @ freckles

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

|-1|=?

OpenStudy (anonymous):

odd

OpenStudy (freckles):

\[F(-x)=\frac{2(-x)}{|-x|}=\frac{2(-x)}{|-1 \cdot x|}=\frac{-2x}{|-1| \cdot |x|}=\frac{-2x}{|x|}=- \frac{2x}{|x|}=-F(x)\] yes odd